OpenStudy (anonymous):

An inductor, L=3mH, a resistor R=8 ohms and a capacitor C=10x10^-6 are connected in series with an AC source whose rms is 25V. find the resonance frequency, and frequencies at which the rms current is 50% of the value of resonance frequency?

OpenStudy (radar):

The formula for resonant frequency is: \[1\over2\pi \sqrt{LC}\]Giving a frequency around 918. Hz inductive reactance XL=2pi F L=17 Ohms Q=Xl/r=17/8=2 bandpass = 918/2=459 at 918 plus and minus 229 Hz the rms current will drop to 70.7% of the value at resonance at 918 Hz